OpenStudy (anonymous):

Which pair has equally likely oucomes? List the letters of the two choices below which have equal pobablities of success,seperate by a comma. A standard deck of cards has 12 face cards and four aces(Aces are not face cards). A) rollinga sum of 11 on two fair sided dice. B) drawing a red nine out of a standard 52 card deck given it''s not a face card or an Ace.C)rolling thm of 10 on two fair six sided dice.D)rollinga sum of 7 on two fair sided dice. E)drawing a six out of a standard 52 card deck given it's not a face card or an Ace.

OpenStudy (anonymous):

A) and B) are equally likely at 2/36=1/13. A) a sum of 11 can be rolled only by 5+6 or 6+5, so there are 2 pairs out of 6*6=36 pairs. 2/36 B) there are 2 red nines. there are 52-16=36 cards that are not a face card or ace. 2/36 If you want to know probabilities for any of the other situations let me know.